# Open3D-VQA: A Benchmark for Embodied Spatial Reasoning with Multimodal Large Language Model in Open Space

### v3

@@ 2.1. Benchmark for Spatial Reasoning
Recent advancements in large language models have demonstrated impressive common-sense reasoning abilities across a wide range of tasks, such as task planning (Shridhar et al., 2020; Wang et al., 2022), navigation (Liu et al., 2023b; Anderson et al., 2018; Krantz et al., 2020), and manipulation. With the integration of multimodal inputs (e.g., images, point clouds), there has been an increasing focus on evaluating the spatial reasoning capabilities of these models. Prior benchmarks focus on four main reasoning categories: (1) relative spatial reasoning (e.g., CLEVR (Johnson et al., 2017), VSR (Liu et al., 2023a)), (2) absolute spatial reasoning (e.g., SpatialVLM (Chen et al., 2024d), SpatialRGPT (Cheng et al., 2024)), (3) situational reasoning involving agent-object relations (e.g., Spatial-MM (Shiri et al., 2024), DriveMLLM (Guo et al., 2024)), and (4) object-centric reasoning (e.g., GPT4Point (Qi et al., 2024), PointLLM (Xu et al., 2025)). However, existing efforts typically cover only a subset of these categories.
@@
2.2. Spatial Reasoning via MLLMs
@@ 4.2. Overall Performance of Baselines
Incorporating point cloud information significantly enhances direction and distance reasoning. All 2D MLLMs perform worse on distance reasoning tasks compared to direction and size reasoning. The average accuracy for distance reasoning across 2D models is only 4.1%, whereas direction and size reasoning reach 33.2% and 40.7%, respectively. The results highlight the challenge current models face in inferring absolute distances. On the other hand, 3D-LLMs achieve comparable performance on distance and direction reasoning, suggesting that point cloud inputs provide models with distance information between objects, thereby strengthening their overall spatial reasoning capability. It is worth noting that although LEO also receives point cloud inputs, its accuracy remains low due to its inability to produce well-formatted responses aligned with the question type, as discussed in Section 4.4.

### v4

@@ 2.1. Benchmark for Spatial Reasoning
Recent advancements in large language models have demonstrated impressive common-sense reasoning abilities across a wide range of tasks, such as task planning (Shridhar et al., 2020; Wang et al., 2022), navigation (Liu et al., 2023b; Anderson et al., 2018; Krantz et al., 2020), and manipulation. With the integration of multimodal inputs (e.g., images, point clouds), there has been an increasing focus on evaluating the spatial reasoning capabilities of these models. Prior benchmarks focus on four main reasoning categories: (1) relative spatial reasoning (e.g., CLEVR (Johnson et al., 2017), VSR (Liu et al., 2023a)), (2) absolute spatial reasoning (e.g., SpatialVLM (Chen et al., 2024d), SpatialRGPT (Cheng et al., 2024)), (3) situational reasoning involving agent-object relations (e.g., Spatial-$M$ (Shiri et al., 2024), DriveMLLM (Guo et al., 2024)), and (4) object-centric reasoning (e.g., GPT4Point (Qi et al., 2024), PointLLM (Xu et al., 2025)). However, existing efforts typically cover only a subset of these categories.
@@
$2.$ Spatial Reasoning via MLLMs
@@ 4.2. Overall Performance of Baselines
Incorporating point cloud information significantly enhances direction and distance reasoning. All 2D MLLMs perform worse on distance reasoning tasks compared to direction and size reasoning. The average accuracy for distance reasoning across 2D models is only 4.1%, whereas direction and size reasoning reach 33.2% and 40.7%, respectively. The results highlight the challenge current models face in inferring absolute distances. On the other hand, 3D-LLMs achieve comparable performance on distance and direction reasoning, suggesting that point cloud inputs provide models with distance information between objects, thereby strengthening their overall spatial reasoning capability. It is worth noting that although LEO also receives point cloud inputs, its accuracy remains low due to its inability to produce well-formatted responses aligned with the question type, as discussed in Section $4.$